ten dry cells were taken from store and voltage tests gave the following results: 1.52,1.53,1.49,1,48,1.47,1.49,1.51,1.50,1.47,1.48 volts. the mean voltage of the cells when stored was 1.51v. assuming the population standard deviation to remain unchanged at 0.02v,


1
Expert's answer
2021-05-14T11:48:20-0400

Given , mean say "x=1.51"


Let "H_o:\mu_o\ge1.51," The mean voltage of the cell greater than 1.51 volts.

and "H_a:\mu_o <1.51" , The mean voltage of the cell is not greater than 1.51 volts.


Population mean, "\mu=\dfrac{14.94}{10}=1.494
,\text{ Population standard deviation }
\sigma=0.02"

Sample size, "n=10"


using t-test for independent sample


Test statistics-


"t=\dfrac{x-\mu}{\frac{\sigma}{\sqrt{n}}}=\dfrac{1.51-1.494}{\frac{0.02}{\sqrt{10}}}=0.9486"



The value of t at 5% level of significance and degree of freedom (n-1) i.e. 9 is 1.895.


Conclusion: As calculated value of t is less than the tabular value of 1.895. So Null hypothesis is accepted. The mean voltage of the cell greater than 1.51 volts.